Find the dimension of a rectangle if the area equals 144 square inches and the length and width Andre in ratio of 2 to 1/2.
Answer by ewatrrr(24785)


Hi
Let x represent the constant of proportionality
A = Lw = 2x(x/2) = x^2 = 144, x = 12 (Tossing out the negative solution for unit measure)
Dimensions are: 24in by 6in
CHECKING our answer***
(24in)(6in) = 144in^2
Wish You the Best in your Studies.